# Modeling Multimodal Social Interactions: New Challenges and Baselines with Densely Aligned Representations

### v1

@@ 4.3 Conversation Context Modeling
| BERT + MViT (Lai et al. [28]) | ✗ | 66.9 | 57.4 |
[…]
@@
| RoBERTa + MViT (Lai et al. [28]) | ✗ | 73.1 | 64.2 |
[…]
@@
| ELECTRA + MViT (Lai et al. [28]) | ✗ | 64.6 | 60.8 |
[…]
@@ 4.4 Aligned Multimodal Fusion
To fuse the aligned visual interaction features and conversation context feature , we first concatenate them along the sequence dimension along with a [AGG] token for feature aggregation. It can be formulated as follows.
Note that positional encoding [49] for transformers is applied to parts. Then, is processed with a multimodal transformer to encode their joint relationships. We leverage an output multimodal feature from the transformed [AGG] token. Finally, a densely aligned multimodal feature is passed through a classification head consisting of an FC layer and softmax to predict the anonymized player identity (e.g., Player#) for the target [MASK]. We optimize the model using cross-entropy loss between the predicted player and ground-truth label .
@@
| BERT + MViT (Lai et al. [28]) | ✗ | 59.8 | 46.4 |
[…]
@@
| RoBERTa + MViT (Lai et al. [28]) | ✗ | 69.5 | 49.1 |
[…]
@@
| ELECTRA + MViT (Lai et al. [28]) | ✗ | 62.1 | 43.8 |
[…]
@@ 5.2 Performance Comparison
Table 2 shows the experimental results for speaking target identification on YouTube and Ego4D datasets with different language models. We compare our proposed baselines with the recent multimodal model [28] (i.e., Language Model + MViT [15]) for social deduction games. In addition, we further adopt DINOv2 [38] which is a powerful versatile visual feature generally used for various downstream tasks. Note that both comparison methods cannot leverage densely aligned language-visual representations. As shown in the table, these methods are not effective in improving upon the language models alone. This reflects that they are not able to figure out who the speaker is and who their gestures are directed at, in correspondence with the language domain. In contrast, our baselines leveraging densely aligned multimodal representations consistently enhance the language models for this task.
Table 3 and 4 show the performance comparison results for the pronoun coreference resolution and mentioned player prediction, respectively. As in the previous task, we conduct experiments with three different language models on two datasets. Competing methods do not perform effectively against the language baselines, but the proposed multimodal baseline performs consistently better than language baselines and the other methods. Our multimodal approach demonstrates the effectiveness of non-verbal visual cues in addressing our introduced social tasks.
@@
| BERT + MViT (Lai et al. [28]) | ✗ | 53.1 | 46.2 |
[…]
@@
| RoBERTa + MViT (Lai et al. [28]) | ✗ | 60.6 | 51.3 |
[…]
@@
| ELECTRA + MViT (Lai et al. [28]) | ✗ | 55.3 | 41.0 |
[…]
@@
[…]
5.3 Effects of Visual Features
We conduct ablation studies on visual feature types to analyze the contribution of each component in our baseline model. Table 5 shows the performance results according to the types of encoded non-verbal cues (i.e., gesture and gaze features) for our social tasks. In these experiments, we adopt the BERT-based baseline for evaluation on YouTube dataset. As shown in the table, the proposed baseline using both gesture and gaze features generally achieves good results. In the case of pronoun coreference resolution, the model utilizing only a gesture feature (i.e., model w/o gaze feature) achieves slightly better performance than our final baseline model. We hypothesize this is because players frequently use hand gestures while referring to other players in the group social deduction game setting.
@@ 5.4 Effects of Conversation Context
Conversational context is typically helpful in understanding a single utterance. To analyze the effects of conversation context for our tasks, we conduct ablation experiments for preceding and following contexts of the target utterance. Table 6 shows the results of different contexts using the BERT-based our baseline on YouTube dataset. A model “w/o preceding and following contexts” is one that uses only the target utterance in terms of language. The last model “w/ both preceding and following contexts” is our proposed baseline. As shown in the table, leveraging both contexts shows the best results for all social tasks. It is noteworthy that the advantage of using the following context is relatively small in the pronoun coreference resolution task compared to the other tasks. It is because the nature of this task is to resolve the reference of pronouns against people that usually appear in the preceding context.
@@ 5.5 Effects of Permutation Learning
To validate the effectiveness of our player permutation learning which shuffles anonymized player identities, we conduct ablation experiments by training models with and without permutation. Table 7 shows the experiment results for three tasks with our BERT-based baseline on YouTube dataset. As shown in the table, the permutation learning approach consistently improves the performances for all tasks, implying it helps the model to learn more generalizable representations of player interactions. Note that we apply this player permutation learning for all comparison methods in Tables 2, 3, and 4 for fair performance comparisons.
@@ 6 Conclusion
We introduce three new challenging tasks in social deduction games: speaking target identification, pronoun coreference resolution, and mentioned player prediction - all of which require understanding the fine-grained verbal and non-verbal dynamics between multiple people. We curate extensive dataset annotations for our new social tasks and further propose a novel multimodal baseline that establishes dense language-visual alignments between spoken utterances and player visual features. This approach enables modeling multi-party social interactions through verbal and non-verbal communication channels simultaneously. Experiments show consistent and considerable performance improvements of our multimodal baselines over other approaches without both modalities, and without dense multimodal alignment. Furthermore, extensive ablation studies are conducted to validate the effectiveness of our baseline components. We will publicly release the benchmarks and code to facilitate further research in this direction.

### v2

@@ 4.3 Conversation Context Modeling
| BERT + MViT [15] (Lai et al. [28]) | ✗ | 66.9 | 57.4 |
| BERT-based Our Baseline | ✓ | 72.7 | 61.9 |
@@
| RoBERTa + MViT [15] (Lai et al. [28]) | ✗ | 73.1 | 64.2 |
| RoBERTa-based Our Baseline | ✓ | 74.5 | 66.5 |
@@
| ELECTRA + MViT [15] (Lai et al. [28]) | ✗ | 64.6 | 60.8 |
| ELECTRA-based Our Baseline | ✓ | 69.6 | 64.8 |
@@ 4.4 Aligned Multimodal Fusion
To fuse the aligned visual interaction features and conversation context feature , we first concatenate them in the sequence dimension along with an [AGG] token for feature aggregation. It can be formulated as follows.
Note that positional encoding [50] for transformers is applied to the parts. Then, is processed with a multimodal transformer to encode their joint relationships. We leverage an output multimodal feature from the [AGG] token. Finally, a densely aligned multimodal feature is passed through a classification head consisting of an FC layer and softmax to predict the anonymized player identity (e.g., Player#) for the target [MASK]. We optimize the model using cross-entropy loss between the predicted player and the ground-truth label .
@@
| BERT + MViT [15] (Lai et al. [28]) | ✗ | 59.8 | 46.4 |
| BERT-based Our Baseline | ✓ | 65.9 | 49.1 |
@@
| RoBERTa + MViT [15] (Lai et al. [28]) | ✗ | 69.5 | 49.1 |
| RoBERTa-based Our Baseline | ✓ | 73.0 | 52.7 |
@@
| ELECTRA + MViT [15] (Lai et al. [28]) | ✗ | 62.1 | 43.8 |
| ELECTRA-based Our Baseline | ✓ | 67.6 | 46.4 |
@@ 5.2 Performance Comparison
Table 2 shows the experimental results for speaking target identification on YouTube and Ego4D datasets with different language models. We compare our proposed baselines with the recent multimodal model [28] (i.e., Language Model + MViT [15]) for social deduction games. In addition, we further adopt DINOv2 [38] which is a powerful versatile visual feature generally used for various downstream tasks. Note that both comparison methods cannot leverage densely aligned language-visual representations. As shown in the table, these methods are not effective in improving upon the language models alone. This reflects that they are not able to figure out who the speaker is and who their gestures are directed at, in correspondence with the language domain. In contrast, our baselines leveraging densely aligned language-visual representations consistently enhance the language models for this task.
Table 3 and 4 show the performance comparison results for pronoun coreference resolution and mentioned player prediction, respectively. We follow a similar experimental setup, evaluating our baselines against the language models and the multimodal methods (i.e., Language Models + DINOv2/MViT) across three different language models and two datasets. The results show that the competing multimodal methods fail to achieve substantial improvements over the language baselines. In contrast, our proposed multimodal baseline consistently outperforms both the language models and the other multimodal methods. Our multimodal approach demonstrates the effectiveness of aligned multimodal cues in addressing these social tasks.
@@
| BERT + MViT [15] (Lai et al. [28]) | ✗ | 53.1 | 46.2 |
| BERT-based Our Baseline | ✓ | 58.8 | 50.0 |
@@
| RoBERTa + MViT [15] (Lai et al. [28]) | ✗ | 60.6 | 51.3 |
| RoBERTa-based Our Baseline | ✓ | 62.5 | 55.1 |
@@
| ELECTRA + MViT [15] (Lai et al. [28]) | ✗ | 55.3 | 41.0 |
| ELECTRA-based Our Baseline | ✓ | 61.0 | 46.2 |
5.3 Effects of Visual Features
We conduct ablation studies on visual feature types to analyze the contribution of each component in our baseline model. Table 5 shows the performance results according to the types of encoded non-verbal cues (i.e., gesture and gaze features) for our social tasks. Our final baseline model encodes speaker kinesics using keypoints related to gaze and gesture, specifically [nose, l-eye, r-eye, l-shoulder, r-shoulder, l-elbow, r-elbow, l-wrist, r-wrist]. “w/o gesture feature” indicates the model utilizing only the head-related keypoints of [nose, l-eye, r-eye] while “w/o gaze feature” employs only the gesture-related keypoints of [l-shoulder, r-shoulder, l-elbow, r-elbow, l-wrist, r-wrist]. We adopt the BERT-based baseline for evaluation on YouTube dataset. As shown in the table, the results show that the gesture features are more dominant compared to the gaze features in our setting. The proposed baselines using both gesture and gaze features generally achieve good performances.
@@ 5.3 Effects of Visual Features
| ✓ | ✓ | 58.8 | |
@@ 5.4 Effects of Conversation Context
Conversational context plays an important role in understanding the meaning and intent behind individual utterances. To analyze the effects of conversation context on our social tasks, we conduct ablation experiments according to the preceding and following contexts of the target utterance. Table 6 presents the results of different context settings using the BERT-based baseline model on YouTube dataset. We compare three variations of the model: one that uses only the target utterance without any additional context, another that incorporates only the preceding context, and our proposed baseline which utilizes both preceding and following contexts. As shown in the table, leveraging both contexts shows the best results for all social tasks. It is noteworthy that the advantage of using the following context is relatively small in the pronoun coreference resolution task compared to the other tasks. This is because the nature of this task is to resolve the reference of pronouns against people that usually appear in the preceding context.
@@ 5.5 Effects of Permutation Learning
To validate the effectiveness of our player permutation learning which shuffles anonymized player identities, we conduct ablation experiments by training models with and without permutation. Table 7 shows the experiment results for three tasks with our BERT-based baseline on YouTube dataset. As shown in the table, the permutation learning approach consistently improves the performances for all tasks, implying it helps the model learn more generalizable representations of player interactions. Note that we apply this player permutation learning to all comparison methods in Tables 2, 3, and 4 for fair performance comparisons.
@@ 6 Conclusion
We introduce three new challenging tasks in social deduction games: speaking target identification, pronoun coreference resolution, and mentioned player prediction - all of which require understanding the fine-grained verbal and non-verbal dynamics between multiple people. We curate extensive dataset annotations for our new social tasks and further propose a novel multimodal baseline that establishes dense language-visual alignments between spoken utterances and player visual features. This approach enables modeling multi-party social interactions through verbal and non-verbal communication channels simultaneously. Experiments show consistent and considerable performance improvements of our multimodal baselines over other approaches without both modalities and without multimodal dense alignment. Furthermore, extensive ablation studies are conducted to validate the effectiveness of our baseline components. We release the benchmarks and source code to facilitate further research in this direction.
Acknowledgement. Portions of this project were supported in part by a gift from Meta.
Appendix A Network Structure Details
Table 8 shows the network structure details of the proposed baseline. The point, kinesics, position, and visual interaction encoders are included in the visual interaction modeling part. The multimodal transformer is used for aligned multimodal fusion. The point, kinesics, and position encoders are implemented with Multilayer Perceptron (MLP) structures, comprising fully connected (FC) layers. The visual interaction encoder and the multimodal transformer are based on typical transformer architectures [50]. For the MLP structures, “MLP Size” denotes the output dimension of each FC layer. ReLU activation is applied between FC layers in MLP structures. For transformers, “Hidden Size” refers to the feature size after passing through the feed-forward network, while “MLP Size” represents the intermediate feature size in the Multi-Head Attention (MHA) mechanism. The channel dimensions and are set to 64 and 512, respectively.
| Network Structures | ||||
| Module | Layers | Hidden Size | MLP Size | Multi-Heads |
| Point Encoder | 3 | – | 64 | – |
[…]
Appendix B Implementation Details
Language Models. We employ ”bert-base-uncased”, “roberta-base”, and “electra-base-discriminator” as pre-trained BERT, RoBERTa, and ELECTRA language models, respectively. We leverage models and weights of them from Hugging Face [52].
Comparison Methods. For comparative analysis, we utilize (Language Model + MViT) and (Language Model + DINOv2). In the case of (Language Model + MViT), we leverage the visual features from the 24-layer multiscale vision transformer (MViT) [15], pre-trained on the Kinetics-400 video dataset, following the approach in [28]. For (Language Model + DINOv2), we use the visual features pooled from a 3-second window of DINOv2 features [38] (interval 0.5s) along the time axis. Both visual features are integrated with the language feature (i.e., conversation context feature) through FC layers for task-specific predictions according to [28].
Appendix C Effects of Conversation Context Length
We conduct experiments to investigate the effects of conversation context length on the performance of each task. Figures 5, 6, and 7 show the validation results for speaking target identification, pronoun coreference resolution, and mentioned player prediction, respectively. When a context length of is employed, the target utterance is concatenated with preceding and following utterances. As shown in the figures, we consistently obtain low performances with the shortest context length of , while achieving fairly good performances with a context length of for all tasks. Note that we adopt as our default setting for the baselines. These evaluations were conducted on YouTube dataset using the BERT-based model.
| Target Task | Conversation Context Length | Accuracy (%) |
| Speaking Target Identification | 66.3 | |
| 70.5 | ||
| 72.7 | ||
| 69.0 | ||
| 69.5 |
| Target Task | Conversation Context Length | Accuracy (%) |
| Pronoun Coreference Resolution | 56.9 | |
| 66.7 | ||
| 65.9 | ||
| 65.9 | ||
| 65.5 |
| Target Task | Conversation Context Length | Accuracy (%) |
| Mentioned Player Prediction | 49.5 | |
| 56.5 | ||
| 58.8 | ||
| 57.3 | ||
| 58.0 |
Appendix D Effects of Video Length
We conduct experiments to investigate the effects of video length on the performance of each task. Figures 8, 9, and 10 show the validation results for speaking target identification, pronoun coreference resolution, and mentioned player prediction, respectively. We achieve fairly good performance with a video length of 3 seconds for all tasks. Note that we adopt 3 seconds as our default setting for the baselines. These evaluations were conducted on YouTube dataset using the BERT-based model.
| Target Task | Video Length | Accuracy (%) |
| Speaking Target Identification | 1 sec | 71.3 |
| 3 sec | 72.7 | |
| 5 sec | 71.5 | |
| 7 sec | 70.8 |
| Target Task | Video Length | Accuracy (%) |
| Pronoun Coreference Resolution | 1 sec | 64.2 |
| 3 sec | 65.9 | |
| 5 sec | 67.4 | |
| 7 sec | 64.6 |
| Target Task | Video Length | Accuracy (%) |
| Mentioned Player Prediction | 1 sec | 56.8 |
| 3 sec | 58.8 | |
| 5 sec | 57.8 | |
| 7 sec | 56.3 |
Appendix E Effects of Player Position Correction
We address scenarios where a player is temporarily undetected, such as being offscreen for a short time. In such cases, we proceed with player position encoding by leveraging the corresponding player position stored in a buffer to correct the missing player position. Table 9 shows the experimental results demonstrating the impact of missing player correction on the performance of all three social tasks. As shown in the table, the position correction contributes to improved performance. These experiments were conducted using the BERT-based model on YouTube dataset.
| Target Task | Player Position Correction | Accuracy (%) |
| Speaking Target Identification | ✗ | 70.1 |
| ✓ | 72.7 | |
| Pronoun Coreference Resolution | ✗ | 65.3 |
| ✓ | 65.9 | |
| Mentioned Player Prediction | ✗ | 58.1 |
| ✓ | 58.8 |
Appendix F Data Domain Generalization
We conduct experiments to validate the generalization capability of our proposed approach across two data domains. To this end, we train our baseline on YouTube dataset and evaluate its performance on Ego4D dataset. Table 10 shows the performance results according to the training data. As shown in the table, the model trained on YouTube data performs well on the Ego4D domain and even achieves better results compared to the model trained on Ego4D for all three social tasks. This improvement can be attributed to the larger amount of training data available in YouTube dataset. The experimental results demonstrate the generalization capability of our approach between different data domains and its potential to work in generalized environments. We adopt the BERT-based model for this experiment.
| Target Task | Test Data | Training Data | Accuracy (%) |
| Speaking Target Identification | Ego4D | Ego4D | 61.9 |
[…]
| Pronoun Coreference Resolution | Ego4D | Ego4D | 49.1 |
[…]
| Mentioned Player Prediction | Ego4D | Ego4D | 50.0 |
[…]
Appendix G Additional Quantitative Results
Utilization of Cropped Visual Features. We conduct experiments using cropped visual image features for visual interaction modeling. This approach is based on our dense alignment framework but utilizes cropped CLIP [42] features instead of keypoint features for the speaker kinesics part (green) in Figure 2 of the main paper. The performances achieved with the cropped CLIP features are 71.0% for Speaking Target Identification, 63.6% for Pronoun Coreference Resolution, and 57.7% for Mentioned Player Prediction. These results are lower compared to our proposed baseline with keypoint features, which achieves 72.7%, 65.9%, and 58.8% for the respective tasks. These evaluations are conducted on YouTube dataset using the BERT-based model.
Measurement of Recall and Precision. In addition to accuracy, we further measure the macro-precision and macro-recall performance for our proposed approach. It is worth noting that in the multi-class setting, accuracy represents both micro-precision and micro-recall. The results show that our approach achieves (macro-precision / macro-recall / accuracy) performances of (74.8% / 74.7% / 72.7%) for Speaking Target Identification, (64.9% / 63.3% / 65.9%) for Pronoun Coreference Resolution, and (61.9% / 60.6% / 58.8%) for Mentioned Player Prediction. These results are obtained using the BERT-based model on YouTube dataset. We could achieve the balanced performances across precision, recall, and accuracy metrics in our environment, considering precision vs recall and macro vs micro aspects.
Appendix H Additional Qualitative Results
Figure 11 illustrates examples where our multimodal baseline model utilizing aligned language and visual cues outperforms the language-only model across the three social tasks. Our baseline with densely aligned multimodal representations enables corrected inferences compared to relying solely on language input. The BERT model is employed for these experiments. When we make inferences for these samples with RoBERTa language-only model, the inference results are #3 for STI, #1 for PCR, and #2 for MPP, which means it fails on 2nd and 3rd samples. These results demonstrate cases where the multimodal model (i.e., BERT) with visual reasoning surpasses the more powerful language-only model (i.e., RoBERTa).
Appendix I Social Deduction Game Details
We leverage two social deduction game datasets: YouTube and Ego4D. Note that the data collections and annotations have been approved by the Institutional Review Board (IRB). YouTube dataset includes the games of One Night Ultimate Werewolf while Ego4D dataset contains the games of One Night Ultimate Werewolf and The Resistance: Avalon. Below are the details for each social deduction game: Werewolf and Avalon.
One Night Ultimate Werewolf. One Night Werewolf is a social deduction game in which players are secretly assigned to one of two primary factions - the villager team or the werewolf team. During the night phase, players close their eyes and characters with special abilities perform actions like swapping cards before opening their eyes again. The night phase may alter players’ roles, though most remain unaware of these changes. Subsequently, players engage in discussion and negotiation to deduce the werewolf’s identity. Werewolves, on their part, strive to conceal their identity and mislead others. At the end, everyone votes on who they believe is most suspicious. If at least one werewolf is eliminated, the village team wins, but if no werewolves are eliminated, the werewolf team wins. We refer One Night Ultimate Werewolf game’s rules on Wikipedia https://en.wikipedia.org/wiki/Ultimate_Werewolf.
The Resistance: Avalon. This game splits players into two groups: the Minions and the Loyal Servants of Arthur. After roles are assigned secretly via card distribution, players commence with a round where each assumes the Leader role in turns. The Leader’s role involves proposing a team for a Quest and all players discuss and vote on approving or rejecting the team assignment. Post the Team Building phase, the designated team decides the Quest’s outcome. The Good Team is restricted to using only the Success card in the Quest phase, whereas the Evil Team has the option to use either the Success or Fail card. The Good Team claims victory upon completing three successful Quests, while the Evil Team wins either by causing three Quests to fail or by correctly identifying the character Merlin among the Good Team. We refer The Resistance: Avalon game’s rules on Wikipedia https://en.wikipedia.org/wiki/The_Resistance_(game).